OpenStudy (anonymous):

What is the y coordinate of the y-intercept of the line that passes through the points (-4, -1) and (4, 5)?

OpenStudy (anonymous):

@cwtan

OpenStudy (anonymous):

my messaging isnt working today

OpenStudy (cwtan):

gradient=5+1/4+4=3/4 y=mx+c 5=3+c c=2